# Generative AI in Transportation Planning: A Survey

### v4

@@ 5.5 Inference with LLMs: Zero-Shot In-Context Learning
Case Study: PNT Estimation Using ZSL. In urban studies, metrics like People Near Transit (PNT) measure the share of the population within proximity (e.g., 500m, 1000m) to transit hubs [271]. To test LLM inference for such tasks, researchers can query LLMs for explicit features such as:
  • Urban planning: How well the city is planned, including the distribution of residential, commercial, and industrial areas, as well as zoning and land use? (0.0: very poor, 10.0: very well planned)
  • Road network quality: How good the road infrastructure is, including the number and condition of roads, highways, and bridges? (0.0: very poor, 10.0: excellent)
  • Public transportation: How extensive and efficient the public transportation system is, including buses, trains, and subways? (0.0: very poor, 10.0: very extensive and efficient)
  • Government policy: How is the government’s investment and development strategy for public transportation? (0.0: very poor, 10.0: very good)
The responses generated by LLMs can be validated by directly correlating their outputs with ground-truth PNT (Proximity to Network Transit) scores [272], where available. This approach provides a robust means of assessing the accuracy of predictions. Notably, this method demonstrates significant improvements compared to directly querying the LLMs for PNT scores without further calibration.
Table 2 summarizes the performance of LLM-generated predictions across varying spatial scales (500m, 1000m, and 1500m) in 85 global cities. The comparison covers three methods: Zero-Shot Learning (ZSL), direct queries, and using average PNT scores as predictions (AVG), and we find that ZSL consistently performs better than traditional methods.
@@
For cities without available transit data, the LLM’s estimations provide a scalable and efficient approximation, which can support decision-making and hypothesis testing.
Comparison with Traditional Methods. Traditional approaches to estimating public transportation accessibility rely on geographic information systems (GIS) and extensive field surveys, which are resource-intensive [273]. ZSL, in contrast, extracts latent knowledge from LLMs using contextual prompts, offering a lightweight alternative for cities with limited data. As demonstrated in [274, 116, 275], LLM-derived features often show strong correlations with real-world transportation metrics, highlighting the potential of ZSL as a viable estimation method.
Strengths and Challenges. ZSL provides significant advantages for transportation planning: It enables rapid scalability across global cities with minimal cost. It can infer latent urban features (e.g., accessibility, traffic density) without training data. Additionally, ZSL introduces a novel user interface paradigm for interacting with transportation models, making advanced modeling tools more accessible and intuitive for practitioners. LLMs can simplify data queries, scenario simulations, and decision-making processes through conversational interfaces, lowering the barrier to entry for complex analysis and fostering broader adoption across the industry. However, LLMs may exhibit uncertainties when tasked with cities for which they lack sufficient knowledge. Such cases often yield generic or inconsistent predictions. Therefore, validating LLM outputs against ground-truth data remains critical for ensuring reliability.
@@ 5.6 Inference with Generative AI: Few-Shot In-Context Learning
Overview of Few-shot Learning. Few-shot in-context learning (FSL) allows large language models (LLMs) to perform complex tasks using only a few examples embedded in a prompt. This capability is particularly beneficial in traffic optimization problems, where annotated data for mixed-autonomy scenarios is limited. In this subsection, we highlight how FSL can be applied in generative AI for traffic coordination using CoMAL (Collaborative Multi-Agent LLMs), a recent framework proposed for mixed-autonomy traffic systems [115].
Designing Effective Few-shot Examples. The effectiveness of few-shot learning heavily relies on the selection of representative examples. For the task of Multi-Agent Traffic Optimization, CoMAL leverages LLM-based multi-agent systems to address the coordination of connected autonomous vehicles (CAVs) in mixed-autonomy traffic, where both autonomous and human-driven vehicles coexist. Few-shot examples are embedded in prompts to enable the LLM to infer strategies for optimizing traffic flow without extensive retraining. In the CoMAL framework, prompts include task descriptions, environment perception, and a few relevant examples of agent roles, driving instructions, and reasoning steps. By analyzing these few-shot examples, the LLM ***For details of LLM version, please visit original paper [115]. coordinates actions between agents, mitigates stop-and-go traffic waves, and enhances overall system efficiency. Few-Shot Coordination in Traffic Merging Scenarios. In the Merge Scenario of CoMAL, vehicles entering from an on-ramp create disturbances in traffic flow on the main highway. Few-shot learning prompts are used to guide the LLM in role allocation and motion planning to reduce congestion and ensure smooth merging behavior. Below is an example of a structured prompt that integrates the traffic task, agent roles, and driving plans.
@@
In this prompt, few-shot examples provide scenarios with varying roles and motion plans. By generalizing from these examples, the LLM infers optimal coordination strategies for traffic merging, balancing acceleration, and deceleration to reduce congestion and shockwaves.
Role of Memory and Collaboration in Few-Shot Learning. In CoMAL, a Memory Module stores previously observed traffic strategies, while a Collaboration Module enables vehicles to share and refine their decisions dynamically. The few-shot examples embedded in prompts are augmented by shared agent observations, ensuring the LLM has context-aware input to reason effectively in diverse traffic scenarios.
Applications in Mixed-Autonomy Traffic. Few-shot in-context learning enables generative AI to perform the following tasks effectively: (1) Role Allocation and Planning: Vehicles autonomously decide on leader or follower roles and adjust speeds for coordinated movement. (2) Traffic Smoothness Optimization: LLM-generated motion instructions mitigate stop-and-go waves by adjusting vehicle headways and speeds dynamically. (3) Generalization to Unseen Scenarios: Few-shot learning allows the system to adapt to novel traffic patterns with minimal examples, enhancing scalability.
Demonstration and Results. Experiments in CoMAL show that LLMs guided by few-shot examples outperform intuitive strategies and achieve smoother traffic flow. As an example in Figure. 5, in the various Merge, Ring, Figure Eight (FE) Scenarios, the system consistently demonstrated improved average vehicle speeds and reduced speed variance compared to human drivers, and even reinforcement learning baselines as discussed in [115].
@@
In conclusion, few-shot in-context learning enables LLMs to efficiently coordinate CAVs in mixed-autonomy traffic systems. By embedding task-specific examples in prompts, LLMs can reason effectively, assign roles, and generate adaptive traffic strategies, demonstrating the power of generative AI in real-time traffic optimization.
Embedding Context in Prompts. Contextual cues are essential in few-shot learning for transportation optimization tasks. In dynamic scenarios like traffic merging or intersection coordination, prompts can be enhanced by including additional information such as traffic density, weather conditions, or vehicle types. These contextual details guide the generative AI model to make more accurate decisions under real-world conditions [115]. For example, specifying that an intersection has “heavy rain” or “high pedestrian activity” prompts the model to account for safety constraints while optimizing traffic flow.
[…]
In this example, contextual cues such as “evening rush hour” and “rainy weather conditions” help the model prioritize safety and efficiency while generating vehicle roles and instructions. These details allow the AI to adapt its recommendations to external factors, improving real-world applicability.
Contextual Adaptation Across Scenarios. Context embedding also enables the model to adapt to diverse traffic situations, such as intersections with: (1) High Pedestrian Activity: Adjust roles to prioritize pedestrian safety over traffic throughput. (2) Emergency Vehicle Presence: Generate instructions to create a clear lane for emergency vehicles. (3) Varying Vehicle Types: Account for differences between connected autonomous vehicles (CAVs) and human-driven vehicles to optimize coordination. By embedding such contextual information in few-shot prompts, generative AI models can reason effectively under complex, dynamic conditions, producing role assignments and strategies that align with real-world transportation priorities [115].
@@ 5.7 Other Techniques Enhancing LLM Inference
In transportation applications, such as intersection signal optimization, Self-Consistency Decoding can stabilize recommendations. For instance, a generative model tasked with adjusting traffic light cycles under heavy congestion may initially produce varying outputs depending on the prompt phrasing. Self-Consistency Decoding generates multiple solutions and selects the one most frequently proposed or aligned with traffic optimization criteria. This approach ensures that the final output reflects a robust, consensus-driven solution that minimizes delays and improves traffic flow.
@@ 5.8 Case Study: LLM’s Ability in OD-Calibration in Terms of Feature Quality
This subsection presents a case study on how LLMs perform in Origin-Destination (OD) Calibration tasks for transportation planning. Specifically, we utilize an O-D calibration dataset from Open Traffic Intelligence (OpenTI) [7], which includes initial OD matrices and observed traffic counts, to evaluate how generative AI models perform when tasked with O-D calibration. The study focuses on two critical aspects: (1) Exploring the few-shot learning solution designs in current LLMs for solving the O-D calibration task, and (2) assessing the quality of model-generated features (observation counts) compared to ground-truth OD matrix’s count feature, an overview of the problem is shown in Figure 6.
Figure 6: The overview of the O-D matrix task in this case study.
@@ 5.8.1 Task Description
In this section, we consider an origin-destination (O-D) matrix $\mathbf{X}$ of dimension $T\times 56\times 56$, where $T$ is the total number of discrete time slices, and $56$ represents the number of blocks in the city. Specifically, $\mathbf{X}$ contains the entries $x_{t,i,j}$ for $t=1,\dots,T$ and $i,j\in\{1,\dots,56\}$, where $x_{t,i,j}\geq 0$ denotes the traffic demand (number of vehicles) from block $i$ to block $j$ during time slice $t$. Once the full matrix $\mathbf{X}$ is specified, a traffic simulator [282] is executed to generate network flows; however, since the observation point is sparse, the effect from the $f\bigl{(}\mathbf{X}\bigr{)}$ is very limited, which brings much difficulty to this task. We denote $O_{t}$ as the ground-truth measurement at the sensor for time slice $t$. Our goal is to calibrate the unknown entries of $\mathbf{X}$ so as to minimize the mean squared error between the simulator-derived flows and the real observations. Concretely, we define the following:
@@
where, $f\bigl{(}x_{t,\cdot,\cdot}\bigr{)}$ denotes the effective flow from simulated traffic that makes differences to observations at the sensor location corresponding to the O-D submatrix (time-slice based O-D) $\{x_{t,i,j}\}_{i,j=1}^{56}$. The discrepancy between the observed and simulated flows is therefore measured by the mean squared error (MSE) at each time step, and the average of these errors over all $T$ time slices forms the objective function in Eq. 1.
@@ 5.8.2 Model Configurations, Computational Resources, and Dataset Selection
We conduct exploration on the open-sourced and widely-used generative AI models Llama 3.1-70B [283]—to simulate O-D calibration scenarios. Experiments are conducted on hardware configurations tailored for large-scale traffic data tasks and are deployed on a server with 8 NVIDIA A100 GPUs, 2 AMD EPYC CPUs, and 2 TB memory.
The dataset used in this case study is from OpenTI [7], a benchmark traffic dataset featuring initial O-D matrices, observed traffic flows, and can generate calibrated O-D outputs for varying network conditions. This dataset captures dynamic traffic patterns across different time intervals, making it ideal for evaluating biases and feature quality in generative O-D calibration tasks.
@@ 5.8.3 Experimental Design and Methodology
Experimental Design. The case study employs two sets of experimental settings, one is the traditional method (Genetic Algorithm), and the other one includes two LLM-based methods.
Method Type 1: Genetic Algorithm for O-D Calibration
To serve as a baseline solution for the O-D calibration problem, we employ a genetic algorithm (GA) using the NSGA2 [284]. The objective is to minimize the mean squared error (MSE) defined in Eq. 1, where each candidate solution represents an estimated O-D matrix $\mathbf{X}$. Given the large search space and non-linear mapping from $\mathbf{X}$ to the observed sensor measurements $O_{t}$, NSGA2 serves as an efficient optimization strategy to iteratively refine O-D matrices that better fit the observed data.
Each candidate O-D matrix $\mathbf{X}$ is represented as a chromosome $\bm{\chi}\in\mathbb{R}^{T\times 56\times 56}$, encoding all $x_{t,i,j}$ values in a flattened vector form. The genetic algorithm initializes a population of such candidate solutions and iteratively evolves them using three key operations: selection, crossover, and mutation. The fitness of each candidate solution is determined by evaluating the simulated sensor flows $f\bigl{(}x_{t,\cdot,\cdot}\bigr{)}$ and computing the corresponding MSE following the Eq. 1.
Formally, the GA optimization in our problem follows these steps:
1.Initialization: A population of $N$ random O-D matrices $\mathbf{X}^{(k)}$, where $k=1,\dots,N$, is generated within predefined bounds (e.g., $x_{t,i,j}\geq 0$).
2.Fitness Evaluation: Each O-D matrix $\mathbf{X}^{(k)}$ is fed into a traffic simulator to generate predicted flows $\widehat{O}_{t}^{(k)}=f\bigl{(}x_{t,\cdot,\cdot}^{(k)}\bigr{)}$. The fitness of each solution is computed by Eq. 1.
3.Selection: Parents are selected based on their fitness scores using NSGA2’s non-dominated sorting and crowding distance mechanisms. Solutions with lower MSE values are more likely to be chosen.
4.Crossover and Mutation: New offspring solutions are generated by recombining parent chromosomes through crossover. Additionally, a mutation operator perturbs selected elements of $\bm{\chi}$ to introduce diversity:
$x_{t,i,j}^{\text{new}}=x_{t,i,j}+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2})$ (2)
where $\sigma$ controls the mutation strength.
5. Survivor Selection: The offspring solutions replace the least fit individuals in the population, ensuring that better solutions persist in subsequent generations.
This iterative process continues until a predefined convergence criterion is met (we set it based on the Interaction steps = 250), with details of setting in Table. 3. The final solution $\mathbf{X}^{*}$ represents the best-calibrated O-D matrix that minimizes the discrepancy between simulated and observed traffic flows.
Table 3: Genetic Algorithm (NSGA2) Hyperparameters for O-D Calibration
@@
Method Type2: LLM-based O-D Calibration Methods
In this section, we have two different designs of LLM-based methods as an exploration. The first one, named VanillaLLM, is the most straightforward design that, we directly pass the current solution with shape $(T\times 56\times 56)$ and the observed link performance, as well as the calculated gap, and the prompt design is as shown in prompt ①. The second design is based on the combination of domain expert knowledge and LLM reasoning, which tries to simplify the problem by leveraging heuristic information to reduce the search space and encouraging LLM’s effective reasoning, named ExpertLLM. The difference between the two LLM-related solvers’ structures is as shown in Figure 7.
@@
Figure 7: The difference between the methodology structure of VanillaLLM and ExpertLLM. As shown in the image, the VanillaLLM directly inputs the flattened O-D matrix to the LLM, together with the MSE value, it provides a vast space to optimize given the dimension of 56 x 56 = 3136 for a single time step, and each value is not bounded, this is a very challenging task for LLM to reason. But ExpertLLM first builds the connection from the O-D matrix to link performance and uses human expert’s domain knowledge on the top-k links searching (which reduces the observation space), then it tracks back to the O-D pairs that contribute to these k links, and LLM is asked to solve a problem on a link basis, which is a much easier task for LLM to infer the adjustment direction, this relaxed the difficulty of the proposed question by leveraging necessary information and combine the analysis with reasoning of LLMs.
The following example prompts illustrate the CoT-based O-D calibration design:
@@
① Prompt for VanillaLLM solving O-D Calibration System Description: We are calibrating a 56×56 Origin-Destination (O-D) matrix for a transportation network. Each entry [i, j] represents trips from origin i to destination j. Known Information: 1. Current O-D matrix (56 lines): {matrix text} 2. Current MSE: {current MSE} Logics: A lower MSE indicates that the O-D matrix better reflects real-world traffic patterns. Task Description: Focus on analyzing patterns and identifying actionable changes to improve matrix accuracy and reduce MSE. Return Constraints: Only reply with 56 lines of a new 56×56 matrix. Example: Adjusted O-D Matrix (for 2x2 shape): $\left[\begin{smallmatrix}320&180\\ 160&340\end{smallmatrix}\right]$. Output: Final O-D matrix with error reduction.

### v5

@@ 4.3.2 Specific Problem Setting: Estimating PNT in Global Cities
Comparison with Traditional Methods
Traditional approaches to estimating public transportation accessibility rely on geographic information systems (GIS) and extensive field surveys, which are resource-intensive [273]. ZSL, in contrast, extracts latent knowledge from LLMs using contextual prompts, offering a lightweight alternative for cities with limited data. As demonstrated in [274, 116, 275], LLM-derived features often show strong correlations with real-world transportation metrics, highlighting the potential of ZSL as a viable estimation method.
Strengths and Challenges
ZSL provides significant advantages for transportation planning: It enables rapid scalability across global cities with minimal cost. It can infer latent urban features (e.g., accessibility, traffic density) without training data. Additionally, ZSL introduces a novel user interface paradigm for interacting with transportation models, making advanced modeling tools more accessible and intuitive for practitioners. LLMs can simplify data queries, scenario simulations, and decision-making processes through conversational interfaces, lowering the barrier to entry for complex analysis and fostering broader adoption across the industry. However, LLMs may exhibit uncertainties when tasked with cities for which they lack sufficient knowledge. Such cases often yield generic or inconsistent predictions. Therefore, validating LLM outputs against ground-truth data remains critical for ensuring reliability.
@@ 4.4 Inference with Generative AI: Few-Shot In-Context Learning
This section presents the general methodology of Few-Shot Learning (FSL) in generative AI and illustrates its application to traffic coordination in mixed-autonomy environments, using the CoMAL framework [115] as an example.
4.4.1 General Methodology: Few-Shot In-Context Learning (FSL)
FSL enables large language models (LLMs) to generalize to new tasks by conditioning on a small number of examples included in a single prompt. It avoids the need for fine-tuning while still leveraging powerful generalization capabilities. Key features of FSL are (1) Minimal supervision: Similar to zero-shot learning, no gradient updates or labeled training datasets are needed. (2) Flexible reasoning: Enables task-specific behavior via few-shot examples and descriptive prompts. (3) Prompt-centered adaptation: The model behavior is guided through structured examples within the input.
4.4.2 Specific Problem Setting: Traffic Coordination in Mixed-Autonomy
In mixed-autonomy scenarios, connected autonomous vehicles (CAVs) and human-driven vehicles share the road. Coordinating these heterogeneous agents in real-time is challenging. The CoMAL framework uses LLMs and FSL to address this coordination challenge in scenarios such as highway merging, roundabouts, and intersections.
Adapting FSL for Traffic Coordination
The effectiveness of few-shot learning heavily relies on the selection of representative examples. For the task of Multi-Agent Traffic Optimization, CoMAL leverages LLM-based multi-agent systems to address the coordination of connected autonomous vehicles (CAVs) in mixed-autonomy traffic, where both autonomous and human-driven vehicles coexist. Few-shot examples are embedded in prompts to enable the LLM to infer strategies for optimizing traffic flow without extensive retraining. In the CoMAL framework, prompts include task descriptions, environment perception, and a few relevant examples of agent roles, driving instructions, and reasoning steps.
1. Few-Shot Prompt Design: Carefully selected coordination examples are embedded into prompts, showing past scenarios of role assignments and motion strategies. In the prompt template titled ‘Few-Shot Prompt for Merging Traffic Coordination’, few-shot examples provide scenarios with varying roles and motion plans. By generalizing from these examples, the LLM infers optimal coordination strategies for traffic merging, balancing acceleration, and deceleration to reduce congestion and shockwaves.
@@ 4.4.2 Specific Problem Setting: Traffic Coordination in Mixed-Autonomy
2. Contextual Reasoning: Context is crucial for tailoring responses to realistic settings. Weather, time of day, and road congestion are embedded in prompts to enable nuanced planning.
[…]
3. Integration with CoMAL Framework: In CoMAL, a Memory Module stores previously observed traffic strategies, while a Collaboration Module enables vehicles to share and refine their decisions dynamically. The few-shot examples embedded in prompts are augmented by shared agent observations, ensuring the LLM has context-aware input to reason effectively in diverse traffic scenarios. FSL prompts are enhanced by this shared context, enabling the LLM to plan cooperatively across agents.
Demonstration and Results
Experiments in CoMAL show that LLMs guided by few-shot examples outperform intuitive strategies and achieve smoother traffic flow. As an example in Figure. 6, in the various Merge, Ring, Figure Eight (FE) Scenarios, the system consistently demonstrated improved average vehicle speeds and reduced speed variance compared to human drivers, and even reinforcement learning baselines as discussed in [115].
@@
Comparison with Traditional Methods
Traditional approaches to traffic coordination often rely on rule-based logic or control-theoretic models. In contrast, FSL with LLMs offers rapid deployment without retraining, adaptability to unseen environments, and seamless integration with real-time, context-rich decision-making.
Strengths and Challenges
FSL supports real-time planning in dynamic traffic, scales to new scenarios with a few examples, and requires no additional labeled datasets.
Dispite its strengths, its performance depends on prompt quality, failure cases can be difficult to interpret or debug, and it needs careful evaluation in safety-critical contexts.
Few-shot in-context learning enables generative AI to coordinate connected autonomous vehicles in complex traffic environments. Through carefully designed prompts and context embedding, LLMs can generate cooperative strategies for traffic merging and other coordination tasks. The approach provides scalable, adaptive, and interpretable decision-making support in mixed-autonomy traffic systems.
@@ 4.5 Other Techniques Enhancing LLM Inference
In transportation applications, such as intersection signal optimization, Self-Consistency Decoding can stabilize recommendations. For instance, a generative model tasked with adjusting traffic light cycles under heavy congestion may initially produce varying outputs depending on the prompt phrasing. Self-Consistency Decoding generates multiple solutions and selects the one most frequently proposed or aligned with traffic optimization criteria. This approach ensures that the final output reflects a robust, consensus-driven solution that minimizes delays and improves traffic.
Figure 7: The overview of the OD matrix task in this case study.
@@ 4.6 Case Study: LLM’s Ability in OD-Calibration in Terms of Feature Quality
This subsection presents a case study on how LLMs perform in Origin-Destination (OD) Calibration tasks for transportation planning. Specifically, we utilize an OD calibration dataset from Open Traffic Intelligence (OpenTI) [7], which includes initial OD matrices and observed traffic counts, to evaluate how generative AI models perform when tasked with OD calibration. The study focuses on two critical aspects: (1) Exploring the few-shot learning solution designs in current LLMs for solving the OD calibration task, and (2) assessing the quality of model-generated features (observation counts) compared to ground-truth OD matrix’s count feature, an overview of the problem is shown in Figure 7.
@@ 4.6.1 Task Description
In this section, we consider an origin-destination (OD) matrix $\mathbf{X}$ of dimension $T\times 56\times 56$, where $T$ is the total number of discrete time slices, and $56$ represents the number of blocks in the city. Specifically, $\mathbf{X}$ contains the entries $x_{t,i,j}$ for $t=1,\dots,T$ and $i,j\in\{1,\dots,56\}$, where $x_{t,i,j}\geq 0$ denotes the traffic demand (number of vehicles) from block $i$ to block $j$ during time slice $t$. Once the full matrix $\mathbf{X}$ is specified, a traffic simulator [282] is executed to generate network flows; however, since the observation point is sparse, the effect from the $f\bigl{(}\mathbf{X}\bigr{)}$ is very limited, which brings much difficulty to this task. We denote $O_{t}$ as the ground-truth measurement at the sensor for time slice $t$. Our goal is to calibrate the unknown entries of $\mathbf{X}$ so as to minimize the mean squared error between the simulator-derived flows and the real observations. Concretely, we define the following:
@@
where, $f\bigl{(}x_{t,\cdot,\cdot}\bigr{)}$ denotes the effective flow from simulated traffic that makes differences to observations at the sensor location corresponding to the OD submatrix (time-slice based OD) $\{x_{t,i,j}\}_{i,j=1}^{56}$. The discrepancy between the observed and simulated flows is therefore measured by the mean squared error (MSE) at each time step, and the average of these errors over all $T$ time slices forms the objective function in Eq. 1.
@@ 4.6.2 Model Configurations, Computational Resources, and Dataset Selection
We conduct exploration on the open-sourced and widely-used generative AI models Llama 3.1-70B [283]—to simulate OD calibration scenarios. Experiments are conducted on hardware configurations tailored for large-scale traffic data tasks and are deployed on a server with 8 NVIDIA A100 GPUs, 2 AMD EPYC CPUs, and 2 TB memory.
The dataset used in this case study is from OpenTI [7], a benchmark traffic dataset featuring initial OD matrices, observed traffic flows, and can generate calibrated OD outputs for varying network conditions. This dataset captures dynamic traffic patterns across different time intervals, making it ideal for evaluating biases and feature quality in generative OD calibration tasks.
@@ 4.6.3 Experimental Design and Methodology
The case study employs two sets of experimental settings, one is the traditional method (Genetic Algorithm), and the other one includes two LLM-based methods.
Table 3: Genetic Algorithm (NSGA2) Hyperparameters for OD Calibration
@@
Method Type 1: Genetic Algorithm for OD Calibration. To serve as a baseline solution for the OD calibration problem, we employ a genetic algorithm (GA) using the NSGA2 [284]. The objective is to minimize the mean squared error (MSE) defined in Eq. 1, where each candidate solution represents an estimated OD matrix $\mathbf{X}$. Given the large search space and non-linear mapping from $\mathbf{X}$ to the observed sensor measurements $O_{t}$, NSGA2 serves as an efficient optimization strategy to iteratively refine OD matrices that better fit the observed data.
@@
Figure 8: The difference between the methodology structure of VanillaLLM and ExpertLLM. As shown in the image, the VanillaLLM directly inputs the flattened OD matrix to the LLM, together with the MSE value, it provides a vast space to optimize given the dimension of 56 x 56 = 3136 for a single time step, and each value is not bounded, this is a very challenging task for LLM to reason. But ExpertLLM first builds the connection from the OD matrix to link performance and uses human expert’s domain knowledge on the top-k links searching (which reduces the observation space), then it tracks back to the OD pairs that contribute to these k links, and LLM is asked to solve a problem on a link basis, which is a much easier task for LLM to infer the adjustment direction, this relaxed the difficulty of the proposed question by leveraging necessary information and combine the analysis with reasoning of LLMs.
Each candidate OD matrix $\mathbf{X}$ is represented as a chromosome $\bm{\chi}\in\mathbb{R}^{T\times 56\times 56}$, encoding all $x_{t,i,j}$ values in a flattened vector form. The genetic algorithm initializes a population of such candidate solutions and iteratively evolves them using three key operations: selection, crossover, and mutation. The fitness of each candidate solution is determined by evaluating the simulated sensor flows $f\bigl{(}x_{t,\cdot,\cdot}\bigr{)}$ and computing the corresponding MSE following the Eq. 1. Formally, the GA optimization in our problem follows these steps: (1) Initialization: A population of $N$ random OD matrices $\mathbf{X}^{(k)}$, where $k=1,\dots,N$, is generated within predefined bounds (e.g., $x_{t,i,j}\geq 0$). (2) Fitness Evaluation: Each OD matrix $\mathbf{X}^{(k)}$ is fed into a traffic simulator to generate predicted flows $\widehat{O}_{t}^{(k)}=f\bigl{(}x_{t,\cdot,\cdot}^{(k)}\bigr{)}$. The fitness of each solution is computed by Eq. 1. (3) Selection: Parents are selected based on their fitness scores using NSGA2’s non-dominated sorting and crowding distance mechanisms. Solutions with lower MSE values are more likely to be chosen. (4) Crossover and Mutation: New offspring solutions are generated by recombining parent chromosomes through crossover. Additionally, a mutation operator perturbs selected elements of $\bm{\chi}$ to introduce diversity: $x_{t,i,j}^{\text{new}}=x_{t,i,j}+\epsilon,\quad\epsilon\sim\mathcal{N}(0,\sigma^{2})$, where $\sigma$ controls the mutation strength. (5) Survivor Selection: The offspring solutions replace the least fit individuals in the population, ensuring that better solutions persist in subsequent generations.
This iterative process continues until a predefined convergence criterion is met (we set it based on the Interaction steps = 250), with details of setting in Table. 3. The final solution $\mathbf{X}^{*}$ represents the best-calibrated OD matrix that minimizes the discrepancy between simulated and observed traffic flows.
Method Type2: LLM-based OD Calibration Methods. In this section, we have two different designs of LLM-based methods as an exploration. The first one, named VanillaLLM, is the most straightforward design that, we directly pass the current solution with shape $(T\times 56\times 56)$ and the observed link performance, as well as the calculated gap, and the prompt design is as shown in prompt ①. The second design is based on the combination of domain expert knowledge and LLM reasoning, which tries to simplify the problem by leveraging heuristic information to reduce the search space and encouraging LLM’s effective reasoning, named ExpertLLM. The difference between the two LLM-related solvers’ structures is as shown in Figure 8.
The following example prompts illustrate the CoT-based OD calibration design:
@@
① Prompt for VanillaLLM solving OD Calibration System Description: We are calibrating a 56×56 Origin-Destination (OD) matrix for a transportation network. Each entry [i, j] represents trips from origin i to destination j. Known Information: 1. Current OD matrix (56 lines): {matrix text} 2. Current MSE: {current MSE} Logics: A lower MSE indicates that the OD matrix better reflects real-world traffic patterns. Task Description: Focus on analyzing patterns and identifying actionable changes to improve matrix accuracy and reduce MSE. Return Constraints: Only reply with 56 lines of a new 56×56 matrix. Example: Adjusted OD Matrix (for 2x2 shape): $\left[\begin{smallmatrix}320&180\\ 160&340\end{smallmatrix}\right]$. Output: Final OD matrix with error reduction.